OpenStudy (anonymous):

A point moves in a straight line so that its distance at time t from a fixed point of the line is 8t - 3t^2. What is the total distance covered by the point between t=1 and t=2?

OpenStudy (anonymous):

\[\int\limits_{1}^{2}8t-3t ^{2} dx\]

OpenStudy (anonymous):

5

OpenStudy (anonymous):

A particle moves in a straight line with velocity v(t) = t^2. How far does the particle move between the times t=1 and t=2? same type of problem, but for some reason I'm getting a strange answer. I integrated to get s(t) and then tried to follow the process above, but...

OpenStudy (anonymous):

\[\int_1^2t^2dt\] yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

why is the answer strange?

OpenStudy (anonymous):

oh... I integrated again because it was asking for position? So I integrated to get s(t) and then again to use the interval?

OpenStudy (anonymous):

i get \[2\tfrac{1}{3}\]

OpenStudy (anonymous):

you want "distance traveled" which is the integral provided the integrand is positive over than interval

OpenStudy (anonymous):

ok. So I don't need to find s(t)?

OpenStudy (anonymous):

you found s(t) when you took the antiderivative of \[t^2\] \[s(t)=\frac{t^3}{3}-1\]

OpenStudy (anonymous):

no that is wrong. \[s(t)=\frac{t^3}{3}-\frac{1}{3}\]

OpenStudy (anonymous):

maybe this is not clear. the derivative of the position function is the velocity and you know the velocity is \[v(t)=t^2\] so the position must be \[s(t)=\frac{t^3}{3}+C\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you also know that you are starting at \[t=1\] so the position there is 0

OpenStudy (anonymous):

(you haven't gone anywhere yet, that is your initial spot)

OpenStudy (zarkon):

Doesn't the first problem give you the distance function

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

Zarkon: yes

OpenStudy (anonymous):

so put \[s(1)=0\] and solve for C you get \[s(1)=\frac{1^3}{3}+C=0\] and so \[C=-\frac{1}{3}\]

OpenStudy (anonymous):

I'm following

OpenStudy (anonymous):

that is one explanation of why when you integrate is is \[F(b)-F(a)\]

OpenStudy (anonymous):

your position function is now known to be \[s(t)=\frac{t^3}{3}-\frac{1}{3}\] so if you want to know where you are at t = 2 just plug it in and get \[s(2)=\frac{2^3}{3}-\frac{1}{3}\]

OpenStudy (anonymous):

which is another way of saying if you want to know how far you have gone compute \[\int_1^2t^2dt\]

OpenStudy (anonymous):

but the problem doesnt actually state that it starts at t=1... that is just the beginning of the interval. Couldn't it already be in motion?

OpenStudy (anonymous):

you are starting at t = 1. think of it this way: if you are traveling at 60 miles per hour it and you want to know how far you travel from 1 pm to 2 pm it doesn't matter when you started your trip, or what mile marker you are at at 1 pm

OpenStudy (anonymous):

oh ok. I understand all of it now. Thank you for your patience

OpenStudy (anonymous):

yw. but maybe we better look at question 1 again because i don't think it is right

OpenStudy (anonymous):

it looks like the first one you are told your position. it says distance is given by \[8t - 3t^2\], not velocity

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so at 1 pm you are at mile marker 5, and at 2 you are at mile marker

OpenStudy (anonymous):

dang thing keep jumping around

OpenStudy (anonymous):

ok this is not trivial. at 1 you are at 5, and at t=4/3 you are at \[5\tfrac{1}{3}\] and then you start heading back and at 2 you are at 4

OpenStudy (anonymous):

so you went 1/3 forward and 2 back for a total distance traveled of \[2\tfrac{1}{3}\]

OpenStudy (anonymous):

that is not one of the answer choices though. If we want to know the total distance why wouldn't we use s(t)?

OpenStudy (anonymous):

what are the choices for #1?

OpenStudy (zarkon):

5/3?

OpenStudy (anonymous):

maybe i messed up. zarkon?

OpenStudy (anonymous):

5/3 is an option. Can you explain your process Zarkon?

OpenStudy (zarkon):

5->51/3=1/3 51/3->4=4/3 1/3+4/3=5/3

OpenStudy (anonymous):

i hope i made a calculation error, not a conceptual one

OpenStudy (zarkon):

i just used your numbers

OpenStudy (anonymous):

oh right, damng

OpenStudy (anonymous):

right from 5 and a third to 4 is 1 and a third. doh

OpenStudy (zarkon):

one can also do this \[f(t)=8t-3t^2\] distance is \[|f(1)-f(4/3)|+|f(4/3)-f(2)|=5/3\]

OpenStudy (anonymous):

not one third forward, 2 back. one third forward and one and one third back

OpenStudy (anonymous):

hope it is clear, even though i screwed up, that we just found the vertex of the parabola, said it was increasing until 4/3 and then decreasing until 2

OpenStudy (anonymous):

I'm really confused where all the thirds are coming from, sorry... would one of you mind starting from the beginning with a concise explanation?

OpenStudy (anonymous):

i messed up so i will let zarkon explain. probably be better

OpenStudy (zarkon):

lol...you were doing fine

OpenStudy (zarkon):

as satellite noticed the position function goes away and then back between the values of t=1 and t=2

OpenStudy (zarkon):

so we need to find when it switches direction

OpenStudy (zarkon):

that is at it's maximum (located at t=4/3) use the vertex formula

OpenStudy (zarkon):

the position at time 1 was f(1)=8(1)-3(1^2)=5 and the position at 4/3 was f(4/3)=5+1/3

OpenStudy (anonymous):

what's the vertex formula? I'm blanking

OpenStudy (zarkon):

for \[ax^2+bx+c\] its \[x=\frac{-b}{2a}\]

OpenStudy (zarkon):

we had \[-3t^2+8t\] \[t=\frac{-8}{2(-3)}=\frac{4}{3}\]

OpenStudy (zarkon):

continuing .... so the distance moving away was (5+1/3)-5=1/3

OpenStudy (anonymous):

but t is going to 2, not 4/3

OpenStudy (anonymous):

4/3 is on the way from 1 to 2

OpenStudy (zarkon):

so now we are ate f(4/3)=5+1/3 and we are moving to f(2)=4 that distance is (5+1/3)-4=1+1/3=4/3

OpenStudy (zarkon):

so the total distance is the 1/3 out plus the 4/3 back 1/3+4/3=5/3

OpenStudy (anonymous):

btw although i generally eschew this explanation, the derivative is the velocity. your derivative is \[-6t+8\] so you can see it is positive (you are going forward) when \[t<\frac{4}{3}\] and negative (you are going back) when \[t>\frac{4}{3}\]

OpenStudy (anonymous):

so you can alway find the vertex by taking the derivative, setting it to zero and solving. of course you will alway get \[-\frac{b}{2a}\] in any case

OpenStudy (anonymous):

turned out to be a bit of a problem didn't it?

OpenStudy (zarkon):

good stuff :)

OpenStudy (anonymous):

\[\checkmark\]

OpenStudy (anonymous):

wow. ok, thanks to both of you. The only thing is I'm not allowed to use a calc for this problem, so how would I necessarily know the graph changes? Or would I go with the velocity method?

OpenStudy (zarkon):

it is a parabola .. it will always change direction...use the vertex formula to find where it changes Or use calculus

OpenStudy (anonymous):

oh ok! So if the vertex was not in the interval, , would I just solve it the way the first guy did it?

OpenStudy (zarkon):

no.... you would evaluate the function at the two values and take the difference

OpenStudy (anonymous):

right ugh sorry, summer makes me forget all my math skills